Lilia is three years older than Rose.Five years ago,the sum of their ages is 41.Find the present age of each.
Found 2 solutions by JBarnum, ikdeep:
Answer by JBarnum(2146)

l=lilla r=rose
given:
l=(3+r)
1-5+r-5=41
l+r=51
_______
(3+r)+r=51
2r=48
r=24
24+3=l
l=27

Answer by ikdeep(226)

lilia's age = 27
and rose = 24
You can see step by step solution from the below URL.
http://www.algebraden.com/solved-problems/find-present-age/29.htm